OpenStudy (anonymous):

prove that if A and B are antisymetric matrix than these are antisymetric too A^T ,A+B,A-B and k*A .please

OpenStudy (anonymous):

antisymetric means:\[A=-A^T,B=-B^T\] So looking at A^T we obtain:\[A^T=-A=-(A^T)^T\]so because:\[A^T=-(A^T)^T\]it is anti symmetric.

OpenStudy (anonymous):

\[A+B=-A^T-B^T=-(A^T+B^T)=-(A+B)^T\]

OpenStudy (anonymous):

The last two will follow the same idea.

OpenStudy (anonymous):

thank you ..i should do the same thing with this too ?:if C is a whatever matrix and C^T *B*C exists and B is antisymetric then C^T *B*C is antisimetryc too

OpenStudy (anonymous):

yeah it will work out the same

OpenStudy (anonymous):

what should I do to prove that if A is a symetric matrix than 2*A^2-3*A+I (unit matrix) is symetric too